# Wider or Deeper? Scaling LLM Inference-Time Compute with Adaptive Branching Tree Search

### v2

@@ 3.2 Adaptive Branching MCTS
Thompson Sampling for Node Selection. In our proposed methods, we employ a Bayesian approach with Thompson sampling for node selection. In standard MCTS, the usual selection policy is based on UCT score [38], but since GEN nodes have no child nodes, UCT scores cannot be assigned. Thompson sampling not only resolves this issue but also enables parallel node expansion. This is particularly beneficial when evaluating node scores is time-consuming, as in the case of MLE-Bench (See Appendix C.1 for MLE-Bench details). Next, we discuss how to utilize Thompson sampling to choose between a GEN node and existing child nodes during the selection step at node $N$. At node $N$, we have $1+n_{\rm child}$ possible actions: choosing a GEN node or one of the $n_{\rm child}$ existing nodes. Since our purpose is to expand the node with score as high as possible, we model the probability distributions of the score of an eventually expanded node at that iteration after choosing actions (i.e. the score of the node $N_{\rm new}$ at Line 5 of Algorithm 1). Thompson sampling proceeds by first sampling scores from the modeled probability distributions for all the actions, and then choosing the action with the highest sampled score. A key question is how to model the expanded node’s score probability distribution for each action. We address this with two strategies: a mixed Bayesian model (AB-MCTS-M) and a node aggregation method (AB-MCTS-A). In both cases, we model the score probability distributions by Bayesian posterior predictives, but with different statistical models.
@@ 4.1 Experimental Setup
Benchmarks. We evaluated AB-MCTS on four diverse benchmarks that require complex problem-solving: LiveCodeBench [14], CodeContest [1], ARC-AGI [17], and MLE-Bench [16]. LiveCodeBench and CodeContest consist of competitive programming problems that demand mathematical and algorithmic reasoning. ARC-AGI involves abstracting a common transformation rule from visual patterns and implementing it as code. MLE-Bench, derived from Kaggle competitions, involves constructing and optimizing machine learning models to achieve high scores based on the evaluation metrics of each competition. For all these benchmarks, LLMs generate Python code to solve each task, and external feedback (e.g., test case results, validation scores) is available to guide the search. More details on the benchmarks can be found in Appendix C.1.
@@
Baselines. We benchmark AB-MCTS against three representative approaches. (1) Repeated Sampling (Best-of-$n$) [1, 3, 39] independently generates up to $n$ candidate solutions from a single LLM prompt, a simple yet competitive baseline for coding tasks. (2) Sequential Refinement [4] iteratively improves each solution by re-prompting the LLM with its own output and feedback. (3) Standard MCTS backs projects such as LATS [9], RAP [15], SWE-Search [10], and RepoUnderstander [11]. As we design a core search algorithm, we group and compare them as standard MCTS. Following the previous work [9, 15], each expansion adds five child nodes, and the search proceeds until it reaches the $2^{7}$ nodes. Hyper-parameters for AB-MCTS are summarized in Appendix C.2.
@@ 4.2 Results
MLE-Bench. Table 2 and Appendix Figure 10 present the performance on three competitions from MLE-Bench using GPT-4o. Since MLE-Bench requires substantial GPU resources for training and evaluating machine learning models, we exclusively used GPT-4o and focused on the baseline methods and AB-MCTS-M (see also Appendix C.1). The best-performing baseline method varies across competitions. This again highlights that different tasks benefit from different exploration-exploitation trade-offs. In contrast, AB-MCTS-M consistently delivers strong performance in these tasks. This consistent success across diverse competitions underscores AB-MCTS-M’s inherent strength in effectively adapting its search strategy to varying problem structures.
@@ 4.3 Analysis
Qualitative Analysis of the Search Trees. Figure 6 and Appendix Figure 11 present example search trees generated by AB-MCTS-M and standard MCTS. These visualizations illustrate more adaptive branching by AB-MCTS-M compared to standard MCTS. This adaptive nature reveals that AB-MCTS-M flexibly balances exploration and exploitation throughout the search process, dynamically allocating budget to explore diverse new candidates ("going wider") and refine promising ones ("going deeper"). Further discussion can be found in Appendix D.3.
@@ A.2.2 Detailed Mixed Model Formulation and Example Code
Here, we adopted the same priors as in our experimental setting, as noted in Appendix C.2:
@@ A.3.2 AB-MCTS-A (Beta) Parameter Update Rules
Appendix B Multi-LLM AB-MCTS
Depending on the task, using more than one LLM for answer search can be advantageous. For example, if LLM A can generate more diverse initial answers and LLM B excels at refinement, building the answer tree with both models is expected to improve performance. This section demonstrates how AB-MCTS can be extended to scenarios in which multiple LLMs are available for answer generation.
B.1 Multiple LLM-Based Answer Generators
Suppose $L$ LLMs are available for answer generation. We denote by $f_{\text{LLM}}^{l}$ the answer generator implemented by the $l$-th LLM, with $l=1,\dots,L$, following the notation introduced in Equation 1. The overall procedure remains identical to single-LLM AB-MCTS (Algorithm 1), except that the expansion phase now includes a generator selection step in which one of the $L$ generators is chosen. This generator selection step is executed on every expansion, using the up-to-date answer tree $T$. The selected generator $f_{\text{LLM}}^{l}$ is then used to expand the current node.
B.2 Generator Selection Algorithm
During this step, each node $N$ in the entire answer tree $T$ is annotated with the index $l$ of the generator that produced it. For every generator, we obtain the set of nodes $\mathcal{N}_{l}\subseteq T,$ containing all nodes it has generated; $\mathcal{N}_{l}$ is empty if the generator has not been selected yet. The scores of the nodes in $\mathcal{N}_{l}$ are used to calculate the posterior distribution of the expected score of a new node that will be generated by the generator $l$. After all $L$ posteriors have been computed, Thompson sampling is applied: a single sample is drawn from each distribution, and the generator with the highest sample is selected for the node expansion. These posteriors are modeled in the same way as the distributions used for node expansion target selection, as detailed below.
B.2.1 Multi-LLM AB-MCTS-M: Generator Selection
Multi-LLM AB-MCTS-M treats the generators as groups within the mixed-effects model defined as
B.2.2 Multi-LLM AB-MCTS-A: Generator Selection
Multi-LLM AB-MCTS-A assigns each $l$-th generator an independent prior–Gaussian (Equations 11-16) or Beta (Equations 17-19)–depending on the score metric, and updates these priors to posteriors using the scores of the nodes in $\mathcal{N}_{l}$. After calculating the posteriors for all $l$, we perform Thompson sampling to choose a generator.
Appendix C Additional Experimental Details
C.1 Tasks and Datasets
@@ C.2 AB-MCTS Parameters
Appendix D Additional Experiments and Analysis
@@
D.1 Results with DeepSeek-V3 on Competitive Programming and ARC-AGI
@@
D.2 Results with GPT-4o on three competitions from MLE-Bench
@@
D.3 Example search trees generated by each methods on MLE-Bench

### v3

@@ 3.2 Adaptive Branching MCTS
Thompson Sampling for Node Selection. In our proposed methods, we employ a Bayesian approach with Thompson sampling for node selection. In standard MCTS, the usual selection policy is based on UCT score [38], but since GEN nodes have no child nodes, UCT scores cannot be assigned. Thompson sampling not only resolves this issue but also enables parallel node expansion. This is particularly beneficial when evaluating node scores is time-consuming, as in the case of MLE-Bench (See Appendix B.1 for MLE-Bench details). Next, we discuss how to utilize Thompson sampling to choose between a GEN node and existing child nodes during the selection step at node $N$. At node $N$, we have $1+n_{\rm child}$ possible actions: choosing a GEN node or one of the $n_{\rm child}$ existing nodes. Since our purpose is to expand the node with score as high as possible, we model the probability distributions of the score of an eventually expanded node at that iteration after choosing actions (i.e. the score of the node $N_{\rm new}$ at Line 5 of Algorithm 1). Thompson sampling proceeds by first sampling scores from the modeled probability distributions for all the actions, and then choosing the action with the highest sampled score. A key question is how to model the expanded node’s score probability distribution for each action. We address this with two strategies: a mixed Bayesian model (AB-MCTS-M) and a node aggregation method (AB-MCTS-A). In both cases, we model the score probability distributions by Bayesian posterior predictives, but with different statistical models.
@@ 4.1 Experimental Setup
Benchmarks. We evaluated AB-MCTS on four diverse benchmarks that require complex problem-solving: LiveCodeBench [14], CodeContest [1], ARC-AGI [17], and MLE-Bench [16]. LiveCodeBench and CodeContest consist of competitive programming problems that demand mathematical and algorithmic reasoning. ARC-AGI involves abstracting a common transformation rule from visual patterns and implementing it as code. MLE-Bench, derived from Kaggle competitions, involves constructing and optimizing machine learning models to achieve high scores based on the evaluation metrics of each competition. For all these benchmarks, LLMs generate Python code to solve each task, and external feedback (e.g., test case results, validation scores) is available to guide the search. More details on the benchmarks can be found in Appendix B.1.
@@
Baselines. We benchmark AB-MCTS against three representative approaches. (1) Repeated Sampling (Best-of-$n$) [1, 3, 39] independently generates up to $n$ candidate solutions from a single LLM prompt, a simple yet competitive baseline for coding tasks. (2) Sequential Refinement [4] iteratively improves each solution by re-prompting the LLM with its own output and feedback. (3) Standard MCTS backs projects such as LATS [9], RAP [15], SWE-Search [10], and RepoUnderstander [11]. As we design a core search algorithm, we group and compare them as standard MCTS. Following the previous work [9, 15], each expansion adds five child nodes, and the search proceeds until it reaches the $2^{7}$ nodes. Hyper-parameters for AB-MCTS are summarized in Appendix B.2.
@@ 4.2 Results
MLE-Bench. Table 2 and Appendix Figure 10 present the performance on three competitions from MLE-Bench using GPT-4o. Since MLE-Bench requires substantial GPU resources for training and evaluating machine learning models, we exclusively used GPT-4o and focused on the baseline methods and AB-MCTS-M (see also Appendix B.1). The best-performing baseline method varies across competitions. This again highlights that different tasks benefit from different exploration-exploitation trade-offs. In contrast, AB-MCTS-M consistently delivers strong performance in these tasks. This consistent success across diverse competitions underscores AB-MCTS-M’s inherent strength in effectively adapting its search strategy to varying problem structures.
@@ 4.3 Analysis
Qualitative Analysis of the Search Trees. Figure 6 and Appendix Figure 11 present example search trees generated by AB-MCTS-M and standard MCTS. These visualizations illustrate more adaptive branching by AB-MCTS-M compared to standard MCTS. This adaptive nature reveals that AB-MCTS-M flexibly balances exploration and exploitation throughout the search process, dynamically allocating budget to explore diverse new candidates ("going wider") and refine promising ones ("going deeper"). Further discussion can be found in Appendix C.3.
@@ A.2.2 Detailed Mixed Model Formulation and Example Code
Here, we adopted the same priors as in our experimental setting, as noted in Appendix B.2:
@@ A.3.2 AB-MCTS-A (Beta) Parameter Update Rules
Appendix B Additional Experimental Details
B.1 Tasks and Datasets
@@ B.2 AB-MCTS Parameters
Appendix C Additional Experiments and Analysis
@@
C.1 Results with DeepSeek-V3 on Competitive Programming and ARC-AGI
@@
C.2 Results with GPT-4o on three competitions from MLE-Bench
C.3 Example search trees generated by each methods on MLE-Bench
@@
Appendix D Multi-LLM AB-MCTS
Depending on the task, using more than one LLM for answer search can be advantageous. For example, if LLM A can generate more diverse initial answers and LLM B excels at refinement, building the answer tree with both models is expected to improve performance. This section demonstrates how AB-MCTS can be extended to scenarios in which multiple LLMs are available for answer generation, and also reports the experimental results for ARC-AGI-2, demonstrating the effectiveness of the proposed method.
D.1 Method
D.1.1 Multiple LLMs as Answer Generators
Suppose $L$ LLMs are available for answer generation. We denote by $f_{\text{LLM}}^{l}$ the answer generator implemented by the $l$-th LLM, with $l=1,\dots,L$, following the notation introduced in Equation 1. The overall procedure is identical to the single-LLM AB-MCTS (Algorithm 1), with the addition of a step to select one of the $L$ available generators for node expansion. This selection occurs at each expansion phase, utilizing the current state of the answer tree $T$. The chosen generator, $f_{\text{LLM}}^{l}$, is then used to expand the selected node. We introduce two distinct algorithms for this generator selection process.
D.1.2 Generator Selection Algorithm I: Single GEN Node
In this algorithm, node selection proceeds identically to AB-MCTS, followed by an additional generator selection step. During this step, each node $N$ in the entire answer tree $T$ is annotated with the index $l$ of the generator that produced it. For every generator, we obtain the set of nodes $\mathcal{N}_{l}\subseteq T,$ containing all nodes it has generated; $\mathcal{N}_{l}$ is empty if the generator has not been selected yet. The scores of the nodes in $\mathcal{N}_{l}$ are used to calculate the posterior distribution of the expected score of a new node that will be generated by the generator $l$. After all $L$ posteriors have been computed, Thompson sampling is applied: a single sample is drawn from each distribution, and the generator with the highest sample is selected for the node expansion. These posteriors are modeled in the same way as the distributions used for node expansion target selection, as detailed below.
Multi-LLM AB-MCTS-M with Generator Selection Algorithm I. In this algorithm, Multi-LLM AB-MCTS-M treats the generators as groups within the mixed-effects model defined as
Multi-LLM AB-MCTS-A with Generator Selection Algorithm I. In this algorithm, Multi-LLM AB-MCTS-A assigns each $l$-th generator an independent prior–Gaussian (Equations 11-16) or Beta (Equations 17-19)–depending on the score metric, and updates these priors to posteriors using the scores of the nodes in $\mathcal{N}_{l}$. After calculating the posteriors for all $l$, we perform Thompson sampling to choose a generator.
D.1.3 Generator Selection Algorithm II: Multiple GEN Nodes
An alternative approach attaches multiple GEN nodes as children to every node in the tree–one for each of the $L$ available generators. At each node $N$ in the expansion target selection process, the following process occurs:
  1. 1.
    The AB-MCTS selection logic is applied independently within the sub-trees associated with each generator. This means that for each generator $l$, we run a selection process considering its associated GEN node and any child nodes previously generated by it from node $N$.
  2. 2.
    Thompson sampling is used to identify the best node (either the GEN node or an existing child node) for each generator $l$.
  3. 3.
    Finally, the scores of the $L$ best nodes selected in the previous step are compared, and the one with the highest overall score is selected.
This algorithm is designed to better capture the local context of the search tree, allowing for the adaptive selection of the most suitable generator at each specific stage of the solution process.
D.2 Experiments
In this section, we evaluate the effectiveness of Multi-LLM AB-MCTS by reporting the results and analysis of ARC-AGI-2 [43]. ARC-AGI-2 is an enhanced benchmark building upon the original ARC-AGI, specifically designed to assess higher-level cognitive abilities of artificial intelligence systems rigorously. It maintains the same fundamental principles as ARC-AGI, emphasizing tasks that require general fluid intelligence rather than extensive prior knowledge or memorization. We selected this benchmark, which is hard to solve even for frontier LLMs (less than 5% success rate [43]), to assess whether we can combine multiple frontier reasoning LLMs to obtain better performance for challenging tasks.
D.2.1 Experimental Setup
In this experiment, we evaluated our approach on the 120 problems comprising the public evaluation set of ARC-AGI-2. For each problem, the generation budget was set to 250. The solution generation and refinement procedures are the same as those of ARC-AGI-1 experiment: The models were instructed to generate the transformation rule as Python code, and the search was guided by a reward signal corresponding to the number of demonstration cases correctly solved by the generated code. For solution generation, we used three frontier reasoning models: Gemini-2.5 Pro (gemini-2.5-pro-preview-05-06[44], o4-mini (o4-mini-2025-04-16[45], and DeepSeek-R1-0528 (deepseek-r1-0528[46]. We set the temperature to 0.6 for all models.
To primarily evaluate the potential of Multi-LLM AB-MCTS, we used the Pass@k metric. This metric measures whether at least one correct solution is found within k attempts. This differs from the official ARC-AGI-2 contest standard, which typically uses a Pass@2 criterion (i.e., one of two submitted final answers must be correct). Evaluating Pass@2 requires an additional selection mechanism to identify promising candidates from the search history. Therefore, this experiment focuses on the search capability itself via Pass@k. Due to the significant API costs of the used frontier reasoning models, we focused on the evaluation of Single-LLM AB-MCTS-A and Multi-LLM AB-MCTS-A (and repeated sampling for o4-mini), since it worked most efficiently in our experiment on ARC-AGI-1. Also, we employed the generator selection algorithm II (see Section D.1.3 for details) for our experiments.
D.2.2 Results
Figure 12: Pass@k (coverage) of ARC-AGI-2 for each generation budget. The methods tested are repeated sampling, AB-MCTS, and Multi-LLM AB-MCTS methods.
The performance of our proposed methods was compared against a repeated sampling, which was the most efficient method for the ARC-AGI-1 experiment, and the result is shown in Fig. 12. As shown, Repeated Sampling using the o4-mini model achieved a 23% Pass@k success rate on the public evaluation set, and the single-model AB-MCTS using o4-mini improved the success rate to 27.5%. The performance advantage of AB-MCTS over Repeated Sampling becomes more evident as the generation budget increases, particularly after approximately 50 budget.
By employing Multi-LLM AB-MCTS, which integrates Gemini-2.5-Pro and DeepSeek-R1-0528, we further improved the performance, ultimately finding correct solutions for over 30% of the problems. Notably, although DeepSeek-R1-0528 exhibited lower individual performance, its integration into the Multi-LLM framework led to an increase in the number of solved problems.
D.2.3 Analysis
Figure 13: Distribution of LLM usage in the 120 ARC-AGI-2 problems with Multi-LLM AB-MCTS-A. Trials are sorted by the maximum reward obtained from demonstration cases (higher reward to the left). Starred trials indicate a correct final solution was found.
The Multi-LLM AB-MCTS framework demonstrated an ability to effectively allocate different LLMs to problems based on their characteristics. As illustrated in Figure 13, the distribution of LLM usage varied across problems. For trials that quickly achieved a high reward from the demonstration cases (left side of the figure), the more proficient model tended to be assigned. Conversely, for trials where obtaining a high reward was more challenging (right side), the models were utilized in a more balanced manner.
Figure 14: An example search tree from a successful trial on ARC-AGI-2 using Multi-LLM AB-MCTS-A. The number in each node indicates the generation step, and the color represents the selected LLM. The yellow node generated the code that correctly solved the test case. This problem was not solved by any single model in isolation.
Figure 15: An illustration of model collaboration. In this example, DeepSeek-R1-0528 refines an incorrect intermediate solution generated by o4-mini (from the problem shown in Figure 14) to produce the final correct solution.
Furthermore, we observed instances where problems unsolvable by any single LLM were solved through the collaboration of multiple models. This suggests a synergistic interaction that transcends simply matching the best model to a problem. Figure 14 and Figure 15 depict a search process where an incorrect solution generated by o4-mini served as a useful hint for DeepSeek-R1-0528 and Gemini-2.5-Pro, which then collaboratively produced the correct solution. This result indicates that Multi-LLM AB-MCTS can facilitate flexible and effective collaboration among heterogeneous frontier LLMs.
D.3 Challenges and Future Work
While our primary evaluation focused on search capability using the Pass@k metric, we conducted a preliminary evaluation based on the Pass@2 criterion for reference. Using a simple rule-based method to select two final answers (prioritizing code with a high reward generated later in the search), the Multi-LLM AB-MCTS achieved a Pass@2 of 19.2%. Although this is a promising result, a significant gap of over 10 percentage points remains compared to the 30% Pass@250 rate. Future work should focus on closing this gap by developing more sophisticated final-answer selection algorithms. Potential directions include building more accurate reward models or integrating an LLM-as-a-Judge for a more nuanced evaluation of candidate solutions.